OpenStudy (anonymous):

Factor and simplify using fundamental identities: csc^3 x - csc^2 x – csc x + 1 The textbook answer is: cot^2 x(csc x – 1). I tried to factor by grouping thus: (csc^3x – csc x) – (csc^2x + 1). But after 1+cot^2 x(cscx – 1) – (csc x +1) I could not make progress.

OpenStudy (lgbasallote):

you did something wrong in your grouping there...

OpenStudy (lgbasallote):

it should be (csc^3 x - csc x) - (csc^2 x - 1) does that make sense?

OpenStudy (lgbasallote):

because it's like you're factoring out -1 in the second group

OpenStudy (lgbasallote):

so if you factor out csc x in the first group you get \[\implies \csc x(\csc^2 x - 1) - (\csc^2 x - 1)\] got it now?

OpenStudy (lgbasallote):

...feel free to ask if there is something you dont understand in what i jus said @SamandAlex

OpenStudy (anonymous):

yes lgbasallote is writehave a look again: .........csc^3x - csc^2x - cscx + 1 ..........(csc^3x - cscx) + (-csc^2x + 1) .........cscx(csc^2 - 1) -(csc^2x-1) remember you need same signs inside the brackets therefore (csc^2x - 1) (cscx - 1) remember this identity:cscx=1/sinx; therefore csc^2x = 1/sin^2x ; substitute this in the brachet you'll get (1/sin^2 - 1) (cscx -1)... in the first brcket u'll get (1 - sin^2x)/sin^2x : look at this numerator remember again sin^2x + cos^2x = 1 ; therefore cos^2x = 1 - sin^2x : substitute this u'll get ( cos^2x/sin^2x) (cscx - 1) ; cos^2x/sin^2x = cot^2x therefore ur text book was right the answer is cot^2x(cscx - 1)

OpenStudy (anonymous):

Thank you. Igbasallote and hubertH. That was very helpful.

OpenStudy (anonymous):

you're welcome